OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

hard for me

OpenStudy (anonymous):

Let number of humans be x and horses be y X + y = 74 2x + 4y =196 so 2x + 2 y = 148 so 2y = 48 and so y = 24 X=,74-24=50 and : 50 humans and 24 horses

OpenStudy (anonymous):

wow great

OpenStudy (anonymous):

wow ur right, good job

OpenStudy (anonymous):

HM = Human HR = Horse HM + HR = 74 2HM + 4HR = 196 (2HM + 4HR) - (2 HM + 2HR) = 196 - 148 2HR = 48 HR = 24 HM + (24) = 74 HM = 74 - 24 HM = 50